OpenStudy (anonymous):

Combinations I must find how many 7 digit registration numbers can start with 1, when the digits available are 0, 1, 2, 3, 4, 5, 6, 7, 9, the registration numbers are issued in ascending order and you cannot have 6 and 9 in the same line,

ganeshie8 (ganeshie8):

looks same q ?

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

you fixed first space wid 1, you're left wid 6 spaces to fill - and you have a constraint that 6 and 9 cannot coexist. Is that right ?

OpenStudy (anonymous):

First, calculate the number of combinations without including the 6th digit.

OpenStudy (anonymous):

yes that is right

ganeshie8 (ganeshie8):

yes thats a good idea, first calculate number of permutations without 6 digit

OpenStudy (anonymous):

k

OpenStudy (anonymous):

c= 7!/(7-6)! c = 5040

ganeshie8 (ganeshie8):

Correct ! thats the total permuations with out digit 6.

OpenStudy (anonymous):

k

ganeshie8 (ganeshie8):

Next, find total permutations with out 9, and with 6 those will be c = 7!/(7-6)! = 5040 but there are some repititions here

OpenStudy (anonymous):

how do u get rid of the repetitions.

ganeshie8 (ganeshie8):

repititions = permutations that doesnt have both 6 and 9 = 6!

OpenStudy (anonymous):

so it is the same?

ganeshie8 (ganeshie8):

so total permutations starting wid 1 w/o 6 and 9 coexisting = 5040 + 5040 - 6!

ganeshie8 (ganeshie8):

this is same as the solution we worked in your previous q, 7! + 6x6!

OpenStudy (anonymous):

here:

ganeshie8 (ganeshie8):

dont ignore the factorial sign \(\color{red}{!}\)

OpenStudy (anonymous):

oh yeah oops

ganeshie8 (ganeshie8):

ha otherwise, it looks good.. replace 6! with 720

OpenStudy (anonymous):

c=9360

ganeshie8 (ganeshie8):

9360 = \(\large \color{red}{\checkmark}\)